OpenStudy (anonymous):

How do I find the first term of a sequence if I am given a random term and the common difference or ratio?

OpenStudy (rina.r):

well try to solve from right and left side both

OpenStudy (anonymous):

How do I know which is right when I get to it?

OpenStudy (rina.r):

then try with division, multiplication, adding, subtracting one of them will give a number of a sequence.

OpenStudy (rina.r):

is ita book question or a worksheet?

OpenStudy (anonymous):

I'm just asking the questions that he said I have to know, so I don't have any practice problems.

OpenStudy (rina.r):

so if u are given a random number and u have to find the 1 number of the s sequence right that what u are asking when u don't know whether it's right or wrong.

OpenStudy (rina.r):

so when u find ur sequence number are same then it will be they right answer.

OpenStudy (anonymous):

Could you type that more clearly, I couldn't understand that.

OpenStudy (rina.r):

ok

OpenStudy (rina.r):

what i am trying to make u understand is that u have to try all these way to get the right number and when u get that number continuously for 2 or 3 time then its a right answer.

OpenStudy (anonymous):

Ok

OpenStudy (rina.r):

now also if u don't understand send me the question i will try to solve it and show u how it works.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

suppose we are given fifth term which is equla to 15 ( which is not frist term) and common difference is 3. let this sequence be in AP Hence \[t_5 = 15 \rightarrow n= 5\] and d= 3 a=? \[t_n = a+(n-1)d \rightarrow t_5 =a+(5-1)3\] \[15 =a+(5-1)3 \rightarrow 15 =a+4 \times 3 \rightarrow 15 =a+12\] i.e. a=15- 12 i.e. a= 3 Hence the first term a= 3 @Ben_Henderson

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

I haven't learned this yet...

OpenStudy (rina.r):

is that an example of find a sequence number?

OpenStudy (rina.r):

*finding